# R2LDM: An Efficient 4D Radar Super-Resolution Framework Leveraging Diffusion Model

### v1

@@ V Conclusion
In this work, we present R2LDM, a novel framework for 4D radar point cloud super-resolution, enhancing autonomous perception in adverse weather. R2LDM employs a latent voxel-based diffusion model (LVDM) to unify radar and LiDAR representations in a shared latent space, effectively capturing 3D structures for improved reconstruction. Built on this foundation, the latent point cloud reconstruction (LPCR) module generates dense, LiDAR-like point clouds while preserving fine-grained details. We introduce a two-stage training strategy, ensuring robust reconstruction across diverse environments. Extensive experiments demonstrate that R2LDM achieves state-of-the-art performance, significantly improving point cloud registration and object detection. These results establish R2LDM as a reliable solution for radar-based autonomous perception, enabling all-weather operation in complex real-world scenarios.

### v2

@@ V Conclusion
In this work, we present R2LDM, a novel framework for 4D radar point cloud super-resolution, enhancing autonomous perception in adverse weather. R2LDM employs a latent voxel-based diffusion model to unify radar and LiDAR in a shared latent space, effectively capturing 3D structures for improved reconstruction. Based on this, the latent point cloud reconstruction module generates dense, LiDAR-like point clouds while preserving fine details. A two-stage training strategy ensures robust reconstruction across diverse environments. Experiments show that R2LDM achieves state-of-the-art performance, significantly improving point cloud registration and object detection, thereby establishing it as a reliable solution for radar-based perception, enabling all-weather operation in complex scenarios. In future work, we will explore how to retain key radar-specific properties, such as velocity and RCS, in enhanced point clouds.